OpenStudy (anonymous):

Can someone help me on this one please? 11 + 4x/15 = 2/3.

Vocaloid (vocaloid):

multiply everything by 15 first

OpenStudy (anonymous):

Sorry, I accidentally put 15 instead of -5, but I did what you said but did -5 and I got\[-55 + \frac{ -20x }{ -5 } = \frac{ -10 }{ 3 }\]

Vocaloid (vocaloid):

hm, ok, let's start over then

Vocaloid (vocaloid):

11 + 4x/(-5) = (2/3)

Vocaloid (vocaloid):

we'll still multiply everything by 15, what do you get?

OpenStudy (anonymous):

\[165 + \frac{ 60 }{ 15 } = \frac{ 30 }{ 3 }\]

OpenStudy (anonymous):

Is this correct?

Vocaloid (vocaloid):

not quite (-4x/5)*15 = ?

OpenStudy (anonymous):

-12?

Vocaloid (vocaloid):

yup

Vocaloid (vocaloid):

(2/3)*15 = ?

OpenStudy (anonymous):

10?